OpenStudy (anonymous):

What is the difference between perfect square trinomials and the difference of two squares?

OpenStudy (mathstudent55):

Perfect square trinomials: \(a^2 + 2ab + b^2 = (a + b)^2 \) \(a^2 - 2ab + b^2 = (a - b)^2 \) Difference of two squares: \(a^2 - b^2 = (a + b)(a - b) \) The square of a binomial results in a trinomial. That is the perfect square trinomial. The product of a sum and a difference results in a the difference of squares which is a binomial.